OpenStudy (anonymous):

the equation of the line passing through the point (-2, 6) and parallel to y = one half x - 8 is ?

OpenStudy (anonymous):

I don't get these at all

OpenStudy (anonymous):

u know the slope is the same if it is parallel, so the equation of the slope is just 1/2= (y-6)/(x+2) slope point form

OpenStudy (anonymous):

is it y = 2x + 5

OpenStudy (anonymous):

y=x/2 + 7

OpenStudy (anonymous):

u needa cross multiply